$SPAD/src/input c06gcf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
y:Matrix SF:=
 [[-0.37168 ,-0.35669 ,-0.31175 ,-0.23702 ,0.00074 ,0.16298 ,1.51370 ]]
n:=ncols y
result:=c06gcf(n,y,-1)

———-

References^{[1] nothing}

4